OpenStudy (anonymous):

What is 2 and 4 over 3 equal to? cube root of 8 square root of 8 cube root of 16 square root of 16

OpenStudy (anonymous):

The question is a fraction, 2 and 3/4

OpenStudy (anonymous):

@Hero @Spacelimbus @austinL

hero (hero):

Do you know what the cube root of 8 is?

OpenStudy (anonymous):

Idk, maybe 2?

hero (hero):

Yes, correct since \(2^3 = 8\)

hero (hero):

What about the square root of 8?

OpenStudy (anonymous):

ummm, 2.8?

hero (hero):

Yes....and how about the cube root of 16?

OpenStudy (anonymous):

2.5

hero (hero):

And we know the square root of 16 is 4.

hero (hero):

Is any of those equal to 2 + 4/3?

hero (hero):

We know that 2 + 4/3 = 2 + 1 + 1/3 = 3 + 1/3

hero (hero):

Wait, was the fraction 2 + 4/3 or 2 + 3/4?

OpenStudy (anonymous):

well, I just want to make sure, but the questions says what is \[2\frac{ 4 }{ 3 }\] equal to?

hero (hero):

Are you sure the answer choices are correct?

hero (hero):

Double check the answer choices

hero (hero):

Sometimes, students post the wrong answer choices for a question without realizing it.

hero (hero):

Make sure the answer choices correspond with the given question.

OpenStudy (anonymous):

yes the answer choices are correct

hero (hero):

I don't know what to tell you then. 2 + 4/3 = 3 + 1/3

hero (hero):

By the way, AND means PLUS

hero (hero):

So \(2 \dfrac{4}{3} = 2 + \dfrac{4}{3}\)

hero (hero):

I prefer to write mixed fractions with the plus to emphasize the correct notation.

OpenStudy (anonymous):

Oh ok, sorry about the and part. Can you help me with another question?

hero (hero):

Close this question and post a new one please.

OpenStudy (anonymous):

ok